# Distributed Sign Momentum with Local Steps for Training Transformers

### v1

@@ C.1 Implementations
Our implementations [^1] https://github.com/shuhuayu/dist-sign-momentum build on nanoGPT [^2] https://github.com/karpathy/nanoGPT and FairScale [^3] https://github.com/facebookresearch/fairscale, where we use nanoGPT for GPT-2 (Radford et al., 2019) modeling, and FairScale for distributed training modules.
@@ C.2 Additional experiment results
Additional validation loss curves. We plot the validation loss curves in Figures 9 and 10 for communication intervals $\tau=24$ and $\tau=36$, respectively. Along with the final validation loss gaps shown in Figure 2, the results indicate improved performance of our proposed sign momentum update compared to the vanilla momentum update in SlowMo.
@@
Figure 10: Validation loss curves for communication interval $\tau=36$.
Details of global AdamW step. In Figure 6, we compare our global sign momentum update with a global update that mimics AdamW, and the corresponding pseudocode is presented in Algorithm 7.
@@
On hyper-parameter fine-tuning. For AdamW, we fine-tune the peak LR and use the recommended values for $(\beta_{1},\beta_{2})$ and weight decay $\lambda$ from Liu et al. (2024). Our results are consistent with those reported in Liu et al. (2024). The same $(\beta_{1},\beta_{2})$ and $\lambda$ values are used for the base optimizer AdamW in both SlowMo and our proposed method. For SlowMo, we fine-tune the local peak LR, global LR, and momentum coefficients. In our proposed method, we adopt the values for $(\beta_{1},\beta_{2})$ and $\lambda$ suggested for Lion in Liu et al. (2024), while fine-tuning only the global LR. Note that our proposed method is more sensitive to communication intervals due to its signed update nature, whereas in SlowMo, the momentum buffer can adapt to the accumulation of gradients from any number of local steps $\tau$, making it less sensitive to communication intervals.

### v2

@@ C.1 Implementations
Our implementations [^1] https://github.com/shuhuayu/dist-sign-momentum build on codebases nanoGPT [^2] https://github.com/karpathy/nanoGPT, FairScale [^3] https://github.com/facebookresearch/fairscale, and Sophia [^4] https://github.com/Liuhong99/Sophia, where we use nanoGPT for GPT-2 (Radford et al., 2019) modeling, Sophia for the optimizer Sophia, and FairScale for distributed training modules. Our algorithm is implemented in the FairScale distributed training modules.
@@ C.2 Additional experiment results
Additional validation loss curves. We plot the validation loss curves in Figures 5 for communication intervals $\tau=24$.
@@
Details of global AdamW step. The detailed description of a global update that mimics AdamW is presented in Algorithm 7.
@@
On hyper-parameter fine-tuning. For AdamW, we fine-tune the peak LR and use the recommended values for $(\beta_{1},\beta_{2})$ and weight decay $\lambda$ from Liu et al. (2024b). Our results are consistent with those reported in Liu et al. (2024b). The same $(\beta_{1},\beta_{2})$ and $\lambda$ values are used for the base optimizer AdamW in both SlowMo and our proposed method. For SlowMo, we fine-tune the local peak LR, global LR, and momentum coefficients. In our proposed method, we adopt the values for $(\beta_{1},\beta_{2})$ and $\lambda$ suggested for Lion in Liu et al. (2024b), while fine-tuning only the global LR. Note that our proposed method is more sensitive to communication intervals due to its signed update nature, whereas in SlowMo, the momentum buffer can adapt to the accumulation of gradients from any number of local steps $\tau$, making it less sensitive to communication intervals.